# The asymptotic dimension of the grand arc graph is infinite

### v2

@@ 4.1.1. Witness projections
Admissible combinatorial models include many familiar graphs, including the curve graph $\mathcal{C}\Omega$, the $1$-skeleton $\mathcal{MC}(\Omega)^{(1)}$ of Masur–Minsky’s marking complex, and the prescribed arc graphs and grand arc graph discussed above.
4.1.2. Combinatorial models on witnesses
Let $\Sigma\subset\Omega$ be a non-annular witness for an admissible combinatorial model $\mathcal{M}$ on $\Omega$. We construct an admissible combinatorial model $\mathcal{M}_{\Sigma}$ on $\Sigma$ for which the projection $\rho_{\Sigma}$ restricts to a Lipschitz map $\mathcal{M}\to\mathcal{M}_{\Sigma}$, along with a Lipschitz coarse section $\iota:\mathcal{M}_{\Sigma}\to\mathcal{M}$. It follows that $\mathcal{M}_{\Sigma}$ quasi-isometrically embeds into $\mathcal{M}$.
Let $V(\mathcal{M}_{\Sigma})=\rho_{\Sigma}(V(\mathcal{M}))$ and let $(a,b)\in E(\mathcal{M}_{\Sigma})$ if and only if $a\neq b$ and there exist $\tilde{a}\in\rho_{\Sigma}^{-1}(a),\tilde{b}\in\rho_{\Sigma}^{-1}(b)$ such that $(\tilde{a},\tilde{b})\in E(\mathcal{M})$. It is immediate that $\rho_{\Sigma}:V(\mathcal{M})\to V(\mathcal{M}_{\Sigma})$ extends to a surjective $1$-Lipschitz map $\rho_{\Sigma}:\mathcal{M}\to\mathcal{M}_{\Sigma}$, hence in particular since $\mathcal{M}$ is connected so is $\mathcal{M}_{\Sigma}$. Likewise, since $\mathcal{M}$ satisfies Definition 4.5(iii), so does $\mathcal{M}_{\Sigma}$ for uniform $L=L_{\Sigma}$. By Corollary 4.4 $\operatorname{PMod}(\Sigma)$ acts naturally on $\mathcal{M}_{\Sigma}$, hence $\mathcal{M}_{\Sigma}$ is admissible.
Fix any $\operatorname{Mod}(\Sigma,\partial\Sigma)$-equivariant section $\iota:V(\mathcal{M}_{\Sigma})\to V(\mathcal{M})$, and let $\tilde{a}=\iota(a)\in\rho_{\Sigma}^{-1}(a)$. We show that $\iota$ is Lipschitz, hence extends to a Lipschitz coarse section $\iota:\mathcal{M}_{\Sigma}\to\mathcal{M}$ for $\rho_{\Sigma}$. Since for any $(a,b)\in E(\mathcal{M}_{\Sigma})$, $i(a,b)\leq L$, there are finitely many $\operatorname{Mod}(\Sigma,\partial\Sigma)$-orbits of edges in $\mathcal{M}_{\Sigma}$. Let
$M=\max_{(a,b)\in E(\mathcal{M}_{\Sigma})/G}d_{\mathcal{M}}(\tilde{a},\tilde{b})$
where $G=\operatorname{Mod}(\Sigma,\partial\Sigma)$. Then $\iota$ is $M$-Lipschitz. We have shown:
@@ 4.1.2. Combinatorial models on witnesses
Let $\Sigma\subset\Omega$ be a non-annular witness for an admissible combinatorial model $\mathcal{M}$ on $\Omega$. There exists an admissible combinatorial model $\mathcal{M}_{\Sigma}$ on $\Sigma$ which quasi-isometrically embeds into $\mathcal{M}$. ∎
@@ 4.2. Asymptotic dimension lower bounds
We first consider the asymptotic dimension of admissible combinatorial models on $\Sigma$, a finite-type surface. Up to deformation retraction, we assume $\Sigma$ is compact.
@@
If $\Sigma$ is a (closed) torus, then any admissible combinatorial model is quasi-isometric to the curve graph, hence a quasi-tree with $\operatorname{asdim}=1$. Otherwise, if $\Sigma$ admits a non-empty admissible combinatorial model (and in particular, a witness subsurface), then $\chi(\Sigma)\leq-1$ and $\Sigma\not\cong\Sigma_{0}^{3}$.
4.2.1. Marking graphs and rank
It suffices to consider admissible marking graphs, in the sense of the following lemma:
Lemma 4.9.
Let $\mathcal{M}$ be an admissible combinatorial model on a compact surface $\Sigma\not\cong\Sigma_{1}$. Then there exists an admissible marking graph $\mathcal{M}^{\prime}$ on $\Sigma$ with an identical witness set and a $\operatorname{PMod}(\Sigma)$-equivariant quasi-isometry $\mathcal{M}\to\mathcal{M}^{\prime}$ which coarsely preserves witness subsurface projection.
For a simple closed curve $a\subset\Sigma$, let $d_{a}(\omega,\omega^{\prime})\mathrel{\mathop{:}}=\operatorname{diam}_{\mathcal{C}(a)}(\pi_{a}\omega\cup\pi_{a}\omega^{\prime})$. Then $d_{a}(\omega,\omega^{\prime})$ is bounded in terms of $i(\omega,\omega^{\prime})$ uniformly in $a$ (see [Wat16, Thm. 2.10], e.g.).
Proof.
Let $\omega=\nu\cup\alpha\in V(\mathcal{M})$. From [Kop23a, §3], we construct a canonical set of locally clean markings $\mu_{\operatorname{base}\omega}$ corresponding to the arc and curve system $\operatorname{base}\omega$ such that
  1. (i)
    for $\mu\in\mu_{\operatorname{base}\omega}$, $\operatorname{base}\nu\subset\mu$ with empty transversals, and
  2. (ii)
    an essential, non-pants subsurface intersects $\mu\in\mu_{\operatorname{base}\omega}$ if and only if it intersects $\operatorname{base}\omega$.
Moreover, there exists $M$ such that for any $\omega,\omega^{\prime}\in V(\mathcal{M})$
  1. (iii)
    $i(\omega,\mu)<M$ for $\mu\in\mu_{\operatorname{base}\omega}$, [^1] This fact is not stated in [Kop23a], but follows from the construction of $\mu_{\operatorname{base}\omega}$. and
  2. (iv)
    if $(\omega,\omega^{\prime})\in E(\mathcal{M})$ then $i(\mu,\mu^{\prime})<M$ for any $\mu\in\mu_{\operatorname{base}\omega},\mu^{\prime}\in\mu_{\operatorname{base}\omega^{\prime}}$.
Obtain the set $\mu_{\omega}$ by adding the transversals in $\nu$ to each $\mu\in\mu_{\operatorname{base}\omega}$. Note that $\mu$ remains locally clean: else, there is a component in $\nu$ bounding an essential, non-pants subsurface disjoint from $\mu$ but not from $\operatorname{base}\omega$, contradicting property (ii). Likewise (iii) still holds for $\mu_{\omega}$.
Let $V(\mathcal{M}^{\prime})=\bigcup_{\omega\in\mathcal{M}}\mu_{\omega}$ and let $(\mu,\mu^{\prime})\in E(\mathcal{M}^{\prime})$ if and only if $\mu\in\mu_{\omega},\mu^{\prime}\in\mu_{\omega^{\prime}}$ for some $(\omega,\omega^{\prime})\in E(\mathcal{M})$. We prove that $\mathcal{M}^{\prime}$ is admissible and $\omega\mapsto\mu_{\omega}$ is the desired (coarse) quasi-isometry. By applying the arguments in [Kop23a] it suffices to verify (ii) and (iv) for the sets $\mu_{\omega}$, replacing $\operatorname{base}\omega$ with $\omega$ and $\operatorname{base}\omega^{\prime}$ with $\omega^{\prime}$, as well as the property
  1. (v)
    if $\mu_{\omega}\cap\mu_{\omega^{\prime}}\neq\varnothing$, then $i(\omega,\omega^{\prime})$ is uniformly bounded.
(ii) for $\mu_{\operatorname{base}\omega}$ implies the same for $\mu_{\omega}$, except for annuli parallel to curves in $\operatorname{base}\nu$; since $\nu\subset\mu\cap\omega$ for $\mu\in\mu_{\omega}$, (ii) holds for $\mu_{\omega}$. Let $\omega=\nu\cup\alpha$ and $\omega^{\prime}=\nu^{\prime}\cup\alpha^{\prime}$. Suppose $(\omega,\omega^{\prime})\in E(\mathcal{M})$ and $\mu\in\mu_{\omega},\mu^{\prime}\in\mu_{\omega^{\prime}}$. Since (iv) holds for $\mu_{\operatorname{base}\omega},\mu_{\operatorname{base}\omega^{\prime}}$, it suffices that $d_{a}(\mu,\mu^{\prime})$ is uniformly bounded for $a\in\nu\cup\nu^{\prime}$; by (ii), if $\mu$ projects to $\mathcal{C}(a)$ then so does $\omega$, and likewise for $\mu^{\prime}$ and $\omega^{\prime}$. $i(\mu,\omega),i(\mu^{\prime},\omega^{\prime})<M$, hence each pair has uniformly close projections if non-empty: $d_{a}(\mu,\mu^{\prime})$ is bounded in terms of $d_{a}(\omega,\omega^{\prime})<L_{\Sigma}$.
Finally, if $\mu\in\mu_{\omega}\cap\mu_{\omega^{\prime}}$, then [Kop23a, §3.2] implies that $i(\operatorname{base}\omega,\operatorname{base}\omega^{\prime})$ is uniformly bounded. For any $a\in\nu\cup\nu^{\prime}$, we show that $d_{a}(\omega,\omega^{\prime})$ is also uniformly bounded. By construction, if $\omega$ or $\omega^{\prime}$ has non-empty projection to $\mathcal{C}(a)$ then so does $\mu$. Since $i(\omega,\mu),i(\omega^{\prime},\mu)<M$, we conclude as above. ∎
@@ 4.2.1. Marking graphs and rank
In particular, by [Kop23a] any admissible marking graph (hence likewise any admissible combinatorial model) $\mathcal{M}$ on a compact surface $\Sigma$ is an asymphoric hierarchically hyperbolic space with respect to subsurface projections to witness curve graphs. Let $\mathscr{X}$ denote the collection of witness subsurfaces for $\mathcal{M}$. Then in particular the rank $\nu$ of $(\mathcal{M},\mathscr{X})$ corresponds to the largest cardinality of a set of pairwise disjoint, connected surfaces in $\mathscr{X}$. Since $(\mathcal{M},\mathscr{X})$ is asymphoric, $\operatorname{asdim}\mathcal{M}\geq\nu$ [BHS21, Thm. 1.15] and $\mathcal{M}$ is $\delta$-hyperbolic if and only if $\nu=1$ [BHS21, Cor. 2.15]. The lower bound here will prove sufficient except when $\nu=1$; we note that an identical bound can be achieved by explicitly constructing quasi-flats.
@@ 4.2.2. The $\delta$-hyperbolic case
Let $\Sigma$ be a genus $g$ compact surface, possibly with boundary. If $\mathcal{M}$ is a (non-empty) $\delta$-hyperbolic admissible combinatorial model on $\Sigma$, then $\operatorname{asdim}\mathcal{M}\geq g-\lceil\frac{1}{2}\chi(\Sigma)\rceil$.
If $\Sigma\cong\Sigma_{1}$, then the claim is immediate by Remark 4.8. Otherwise, we may assume $\mathcal{M}$ is an admissible marking graph by Lemma 4.9. Let $\mathscr{X}^{\mathcal{M}}$ denote the collection of witness subsurfaces for $\mathcal{M}$. For any $\mathcal{M}^{\prime}$ an admissible marking graph on $\Sigma$ with $\mathscr{X}^{\mathcal{M}}\supset\mathscr{X}^{\mathcal{M}^{\prime}}$, there exists a functorial canonical coarse surjection $\iota:\mathcal{M}\to\mathcal{M}^{\prime}$ such that $\pi_{W}\circ\iota$ is uniformly coarsely $\pi_{W}$ for any $W\in\mathscr{X}^{\mathcal{M}^{\prime}}$ [Kop23a, §2.1]. In particular, $\mathscr{X}^{\mathcal{MC}^{(1)}(\Sigma)}$ is every essential, non-peripheral subsurface in $\Sigma$ and $\mathscr{X}^{\mathcal{C}\Sigma}=\{\Sigma\}$, hence we have canonical maps $\mathcal{MC}^{(1)}(\Sigma)\to\mathcal{M}\to\mathcal{C}\Sigma$.
@@
Let $\mathcal{M},\mathcal{M}^{\prime}$ be admissible marking graphs on $\Sigma$, a compact surface, such that $\mathscr{X}^{\mathcal{M}}\supset\mathscr{X}^{\mathcal{M}^{\prime}}$, and let $\iota:\mathcal{M}\to\mathcal{M}^{\prime}$ be the canonical coarse surjection. If $\mathcal{M}$ is $\delta$-hyperbolic, then $\iota$ is coarsely alignment-preserving.
@@
Suppose that $\mathcal{M}$ is a $\delta$-hyperbolic admissible marking graph on a compact surface $\Sigma$ with genus $g$. By Lemma 4.11, the canonical map $\iota:\mathcal{M}\to\mathcal{C}\Sigma$ is coarsely alignment preserving, hence by Theorem 1.12 $\partial\mathcal{C}\Sigma$ embeds into $\partial\mathcal{M}$. To prove Theorem 4.10 it suffices to find a compact subspace $Z\subset\partial\mathcal{C}\Sigma$ such that $\dim Z\geq n\mathrel{\mathop{:}}=g-1-\lceil\frac{1}{2}\chi(\Sigma)\rceil$, since by Proposition 2.5 $\dim Z+1\leq\operatorname{asdim}\mathcal{M}$. Recall that $\partial\mathcal{C}\Sigma\cong\mathcal{EL}(\Sigma)$.
@@
When $\Sigma$ is a sphere with four boundary components, Theorem 4.10 is vacuously true. Otherwise, from Theorem 2.4 and the universal embedding property of Nöbeling spaces, we obtain the desired subspace $Z\subset\operatorname{\mathcal{EL}}(S)\subset\partial\mathcal{C}\Sigma$ and Theorem 4.10 follows. ///\!\!// /
@@ 4.2.3. Lower bounds for infinite-type surfaces
Given an admissible combinatorial model $\mathcal{M}$ on an infinite-type surface $\Omega$, let $w_{\mathcal{M}}\in\mathbb{N}\cup\{\infty\}$ denote the least upper bound on cardinalities for a set of pairwise-disjoint connected witnesses for $\mathcal{M}$. We consider two cases:
@@
Let $\mathcal{M}$ be an admissible combinatorial model on an infinite-type surface $\Omega$. Then $\operatorname{asdim}\mathcal{M}=\infty$. ∎
Theorem 1.3 follows from Theorems 4.10 and 4.13 in the special case of admissible multiarc and curve graphs. ///\!\!// /

### v3

@@ 4.2.1. Witness projections
Witness-cocompact arc and curve models include many graphs of contemporary interest on infinte-type surfaces, including the ray graph, the omnipresent arc graph, and the grand arc graph.
4.2.2. Arc and curve models on witnesses
Let $\Sigma\subset\Omega$ be a witness for a witness-cocompact arc and curve model $\mathcal{M}$ on $\Omega$. We construct a cocompact arc and curve model $\mathcal{M}_{\Sigma}$ on $\Sigma$ for which the projection $\rho_{\Sigma}$ restricts to a Lipschitz map $\mathcal{M}\to\mathcal{M}_{\Sigma}$, along with a Lipschitz coarse section $\iota:\mathcal{M}_{\Sigma}\to\mathcal{M}$. It follows that $\mathcal{M}_{\Sigma}$ quasi-isometrically embeds into $\mathcal{M}$.
Let $V(\mathcal{M}_{\Sigma})=\rho_{\Sigma}(V(\mathcal{M}))$ and let $(a,b)\in E(\mathcal{M}_{\Sigma})$ if and only if $a\neq b$ and there exist $\tilde{a}\in\rho_{\Sigma}^{-1}(a),\tilde{b}\in\rho_{\Sigma}^{-1}(b)$ such that $(\tilde{a},\tilde{b})\in E(\mathcal{M})$. It is immediate that $\rho_{\Sigma}:V(\mathcal{M})\to V(\mathcal{M}_{\Sigma})$ extends to a surjective $1$-Lipschitz map $\rho_{\Sigma}:\mathcal{M}\to\mathcal{M}_{\Sigma}$, hence in particular since $\mathcal{M}$ is connected so is $\mathcal{M}_{\Sigma}$. Likewise, since $\mathcal{M}$ satisfies Definition 4.16(iii), so does $\mathcal{M}_{\Sigma}$ for uniform $L=L_{\Sigma}$. By Corollary 4.15 $\operatorname{PMap}(\Sigma)$ acts naturally on $\mathcal{M}_{\Sigma}$, hence $\mathcal{M}_{\Sigma}$ is a cocompact arc and curve model on $\Sigma$.
Fix any $\operatorname{Map}(\Sigma,\partial\Sigma)$-equivariant section $\iota:V(\mathcal{M}_{\Sigma})\to V(\mathcal{M})$, and let $\tilde{a}=\iota(a)\in\rho_{\Sigma}^{-1}(a)$. We show that $\iota$ is Lipschitz, hence extends to a Lipschitz coarse section $\iota:\mathcal{M}_{\Sigma}\to\mathcal{M}$ for $\rho_{\Sigma}$. Since for any $(a,b)\in E(\mathcal{M}_{\Sigma})$, $i(a,b)\leq L$, there are finitely many $\operatorname{Map}(\Sigma,\partial\Sigma)$-orbits of edges in $\mathcal{M}_{\Sigma}$. Let
[…]
where $G=\operatorname{Map}(\Sigma,\partial\Sigma)$. Then $\iota$ is $M$-Lipschitz. We have shown:
@@ 4.2.2. Arc and curve models on witnesses
Let $\Sigma\subset\Omega$ be a witness for a witness-cocompact arc and curve model $\mathcal{M}$ on $\Omega$. There exists a cocompact arc and curve model $\mathcal{M}_{\Sigma}$ on $\Sigma$ which quasi-isometrically embeds into $\mathcal{M}$. ∎
Remark 4.19.
A connected arc and curve model $\mathcal{M}$ on $\Omega$ with a natural action of $\operatorname{PMap}_{c}(\Omega)$ is witness-cocompact if and only if it has a witness and the projection $\mathcal{M}_{\Delta}$ is cocompact for every witness $\Delta\subset\Omega$.
@@ 4.3. Asymptotic dimension lower bounds
We first consider the asymptotic dimension of cocompact arc and curve models on $\Sigma$, a finite-type non-pants hyperbolic surface. Up to deformation retraction, we assume $\Sigma$ is compact.
@@
If $\Sigma$ is a (closed) torus, then any cocompact arc and curve model is quasi-isometric to the curve graph, hence a quasi-tree with $\operatorname{asdim}=1$. Otherwise, if $\Sigma$ admits a non-empty cocompact arc and curve model (and in particular, a witness subsurface), then $\chi(\Sigma)\leq-1$ and $\Sigma\not\cong\Sigma_{0}^{3}$.
@@
Let $\mathcal{M}$ be a cocompact arc and curve model on $\Sigma$; recall that $(\mathcal{M},\mathscr{X}^{\mathcal{M}})$ is an asymphoric hierarchically hyperbolic space by Theorem 4.12. Then in particular the rank $\nu$ of $(\mathcal{M},\mathscr{X}^{\mathcal{M}})$ corresponds to the largest cardinality of a set of pairwise disjoint, connected witnesses in $\mathscr{X}^{\mathcal{M}}$. Since $(\mathcal{M},\mathscr{X}^{\mathcal{M}})$ is asymphoric, $\operatorname{asdim}\mathcal{M}\geq\nu$ [BHS21, Thm. 1.15] and $\mathcal{M}$ is $\delta$-hyperbolic if and only if $\nu=1$ [BHS21, Cor. 2.15]. The lower bound here will prove sufficient except when $\nu=1$; we note that an identical bound can be achieved by explicitly constructing quasi-flats.
@@ 4.3.1. The $\delta$-hyperbolic case
Let $\Sigma$ be a genus $g$ compact surface, possibly with boundary. If $\mathcal{M}$ is a (non-empty) $\delta$-hyperbolic cocompact arc and curve model on $\Sigma$, then $\operatorname{asdim}\mathcal{M}\geq g-\lceil\frac{1}{2}\chi(\Sigma)\rceil$.
If $\Sigma\cong\Sigma_{1}$, then the claim is immediate by Remark 4.20. Otherwise, we may assume $\mathcal{M}$ is a cocompact marking graph by Proposition 4.3. For any $\mathcal{M}^{\prime}$ a cocompact marking graph on $\Sigma$ with $\mathscr{X}^{\mathcal{M}}\supset\mathscr{X}^{\mathcal{M}^{\prime}}$, there exists a functorial canonical coarse surjection $\iota:\mathcal{M}\to\mathcal{M}^{\prime}$ such that $\pi_{W}\circ\iota$ is uniformly coarsely $\pi_{W}$ for any $W\in\mathscr{X}^{\mathcal{M}^{\prime}}$ [Kop23a, §2.1]. In particular, $\mathscr{X}^{\mathcal{MC}(\Sigma)}$ is every essential, non-peripheral subsurface in $\Sigma$ and $\mathscr{X}^{\mathcal{C}\Sigma}=\{\Sigma\}$, hence we have canonical maps $\mathcal{MC}(\Sigma)\to\mathcal{M}\to\mathcal{C}\Sigma$.
@@
Let $\mathcal{M},\mathcal{M}^{\prime}$ be cocompact marking graphs on $\Sigma$, a compact surface, such that $\mathscr{X}^{\mathcal{M}}\supset\mathscr{X}^{\mathcal{M}^{\prime}}$, and let $\iota:\mathcal{M}\to\mathcal{M}^{\prime}$ be the canonical coarse surjection. If $\mathcal{M}$ is $\delta$-hyperbolic, then $\iota$ is coarsely alignment-preserving.
@@
Suppose that $\mathcal{M}$ is a $\delta$-hyperbolic cocompact marking graph on a compact surface $\Sigma$ with genus $g$. By Lemma 4.22, the canonical map $\iota:\mathcal{M}\to\mathcal{C}\Sigma$ is coarsely alignment preserving, hence by Theorem 1.13 $\partial\mathcal{C}\Sigma$ embeds into $\partial\mathcal{M}$. To prove Theorem 4.21 it suffices to find a compact subspace $Z\subset\partial\mathcal{C}\Sigma$ such that $\dim Z\geq n\mathrel{\mathop{:}}=g-1-\lceil\frac{1}{2}\chi(\Sigma)\rceil$, since by Proposition 2.5 $\dim Z+1\leq\operatorname{asdim}\mathcal{M}$. Recall that $\partial\mathcal{C}\Sigma\cong\mathcal{EL}(\Sigma)$.
@@
When $\Sigma$ is a sphere with four boundary components, Theorem 4.21 is vacuously true. Otherwise, from Theorem 2.4 and the universal embedding property of Nöbeling spaces, we obtain the desired subspace $Z\subset\operatorname{\mathcal{EL}}(S)\subset\partial\mathcal{C}\Sigma$ and Theorem 4.21 follows. ///\!\!/
@@ 4.3.2. Lower bounds for infinite-type surfaces
Given a witness-cocompact arc and curve model $\mathcal{M}$ on an infinite-type surface $\Omega$, let $w_{\mathcal{M}}\in\mathbb{N}\cup\{\infty\}$ denote the least upper bound on cardinalities for a set of pairwise-disjoint connected witnesses for $\mathcal{M}$. We consider two cases:
@@
Let $\mathcal{M}$ be a witness-cocompact arc and curve model on an infinite-type surface $\Omega$. Then $\operatorname{asdim}\mathcal{M}=\infty$. ∎
Theorem 1.3 follows from Theorems 4.21 and 4.24. ///\!\!/